# Causally Reliable Concept Bottleneck Models

### v2

@@ Appendix F Experimental details
Python code and instructions for reproducing results across all datasets and methods are available within the code provided alongside the submission as supplementary material. We detail below the configurations and hyperparameters used for models’ instantiation and training. For a fair comparison across concept-based models, we standardize the concept encoder to a single-hidden-layer MLP. All models are trained using the Adam optimizer (Kingma & Ba, 2015) for a maximum of 500 epochs, with early stopping based on a 30-epoch patience. We employ LeakyReLU as the activation function throughout.
The batch size is set to 512 for most datasets, with the exception of Siim-pneumothorax and SCBM, where it is reduced to 128 due to memory constraints. Following Koh et al. (2020), we regularize the task loss to encourage concept learning, using a weighted sum of task and concept losses:
@@
Additionally, we apply random training-time interventions as proposed by Zarlenga et al. (2022), with an intervention probability of 0.25. Regarding SCBM, we used the authors’ implementation at https://github.com/mvandenhi/SCBM. More precisely, we implemented the global variation using the configuration proposed by the authors.
@@ F.1 Metric: custom Structural Hamming Distance
$i-j$ / $1/4$
$i-j$ $i\rightarrow j$ $1/4$
$i\rightarrow j$ $i-j$ $1/5$
Table 5: SHD penalties for various discrepancies between ground truth and predicted edges. ’$i\rightarrow j$’: oriented edge, ’$i-j$’: unoriented edge, ’///’: no edge
@@ G.1 Label accuracy
Table 6: Label accuracy (%). Task concepts are as follows: dysp (Asia), Akt (Sachs), BP (Alarm), PropCost (Insurance), R5Fcst (Hailfinder), parity (cMNIST), mouth slightly open (CelebA), Survival (CUBC), pneumothorax (Pneumoth.). Uncertainties represent $2$ sample mean $\sigma$ (5 seeds).
@@ G.2 Task accuracy using true graph
Table 7: Task accuracy (%) using the true and predicted graph. Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder). Uncertainties represent $2$ sample mean $\sigma$ (5 seeds).
@@
C2BM True graph $71.0_{\pm 2.3}$ $65.0_{\pm 1.3}$ $66.4_{\pm 1.8}$ $62.1_{\pm 1.5}$ $73.0_{\pm 1.6}$
Predicted graph $71.4_{\pm 1.7}$ $65.3_{\pm 1.1}$ $66.4_{\pm 1.5}$ $62.5_{\pm 1.4}$ $74.1_{\pm 1.8}$
@@
G.3 Effect of single-concept interventions
G.4 Ablation study on Causal Discovery
Table 8 evaluates the sensitivity of the causal discovery component in our pipeline to the choice of method for constructing a causal graph. The results are evaluated on all the bnlearn datasets, for which the true causal graphs are known.
We focus on comparing the Greedy Equivalence Search (GES) algorithm, our method of choice, with different widely-used causal discovery algorithms that recover a CPDAG. These include the PC algorithm and Fast GES, described below, each evaluated both with and without refinement via the retrieval-augmented generation (RAG)-enhanced language model (LLM) employed in our study. Furthermore, we compare these methods against the use of the LLM alone, as well as our retrieval-augmented LLM (LLM + RAG) applied directly for causal graph recovery. Specifically, the LLM is prompted with the causal prompt shown in App. C.3, with additional retrieved context appended to the query in the LLM+RAG setting
@@ G.4 Ablation study on Causal Discovery
  • GES (score-based): The algorithm selected in our study. A detailed description is provided in App. C.
  • PC Algorithm (independence-based): A classical independence-based method that first estimates the undirected structure of the causal relations through a sequence of conditional independence tests, then orients edges using a set of predefined orientation rules (Colombo et al., 2014). In our implementation, we use the PC algorithm provided by the causallearn library, with a chi-squared independence test and a significance level of 0.05.
  • Fast GES (FGES) (score-based): A computationally efficient variant of GES that improves performance by storing intermediate evaluations and parallelizing expensive operations, enabling its application to large datasets (Andrews et al., 2019; Ramsey et al., 2017). In our implementation, we use the FGES algorithm provided by the py-tetrad library, employing the Bayesian Dirichlet Equivalent Uniform (BDeu) score for structure evaluation (Heckerman et al., 1995).
Table 8: Structural Hamming distance and mistaken edge ratio between the true and the learned causal graphs for each tested method.
@@
Mistaken Max errors 28 55 351 666 1540
edges Random Classifier 9.33 18.33 117 222 513.33
ratio LLM 13 20 351 543 107
LLM + RAG 12 12 226 Time limit Time limit
PC 6 10 26 13 45
FGES 3 17 23 14 39
GES 3 17 19 13 22
PC + LLM (+ RAG) 6 9 22 9 43
FGES + LLM (+ RAG) 1 8 15 11 36
GES + LLM (+ RAG) 1 7 18 10 22
@@
As shown in the results, GES refined with the RAG-augmented LLM is the overall best-performing method in the majority of cases, both in terms of structural Hamming distance and mistaken edge ratio. In instances where it is not the best, it still consistently ranks among the top methods. Notably, both the standard causal discovery method (GES) and the use of the RAG-augmented LLM contribute positively to performance. Due to the substantial computational time required by the LLM+RAG-based causal discovery approach, we excluded experiments on datasets for which execution exceeded 24 hours.
G.5 Ablation study on LLM type
Tab. 9 presents the sensitivity of the causal discovery component when different LLMs are employed. In this experiment, both LLMs utilize the context provided by the RAG to respond to the causal queries. GPT-4o and GPT-4o-mini are both LLMs developed by OpenAI, with the former being a more powerful, computationally intensive model compared to the latter. The GPT-4o-mini is more efficient, delivering better performance than older models such as GPT-3.5-turbo on most benchmarks, while significantly reducing computational costs compared to GPT-4o. Since both models are not open-source, additional details are not available.
Table 9: Structural Hamming distance and number of mistaken edges between true and learned DAG when using either GPT-4o or GPT-4o-mini.
Metric LLM Sachs Alarm
Hamming GPT-4o-mini 2.16 5.0
GPT-4o 1.8 5.0
Mistaken GPT-4o-mini 8 (17) 10 (46)
edges ratio GPT-4o 7 (17) 10 (46)
Analyzing the results in Tab. 9, we observe that on the Alarm dataset, both LLMs achieve identical scores for both metrics. However, the smaller GPT-4o-mini model performs less effectively on the Sachs dataset, as it incorrectly identifies an additional edge, resulting in a less accurate graph and a higher Hamming index. Consequently, we conclude that the performance of the causal discovery component is influenced by the choice of LLM, although the performance differences between the two models are minimal for the datasets tested.
G.6 Ablation study on RAG
@@
Mistaken No context 12 (17) 10 (46)
edges ratio RAG context 7 (17) 10 (46)
@@
G.7 Decomposing interventional accuracy
@@
Figure 12: Task accuracy improvement (%) in predicting downstream variables after intervening on groups of concepts up to progressively deeper levels in the graph hierarchy. The metric is averaged across all downstream variables (Left), the task only (Middle), and all concepts (Right) Total label accuracy. Uncertainties represent $1.96$ sample mean $\sigma$ (5 seeds).

### v3

@@ Appendix F Experimental details
Python code and instructions for reproducing results across all datasets and methods are available within the code provided alongside the submission as supplementary material. We detail below the configurations and hyperparameters used for models’ instantiation and training. For a fair comparison across concept-based models, we standardize the concept encoder to a single-hidden-layer MLP. All models are trained using the Adam optimizer (kingma2014adam) for a maximum of 500 epochs, with early stopping based on a 30-epoch patience. We employ LeakyReLU as the activation function throughout.
The batch size is set to 512 for most datasets, with the exception of Siim-pneumothorax and SCBM, where it is reduced to 128 due to memory constraints. Following koh2020, we regularize the task loss to encourage concept learning, using a weighted sum of task and concept losses:
@@
Additionally, we apply random training-time interventions as proposed by zarlenga2022concept, with an intervention probability of 0.25. Regarding SCBM, we used the authors’ implementation at https://github.com/mvandenhi/SCBM. More precisely, we implemented the global variation using the configuration proposed by the authors.
@@ F.1 Metric: custom Structural Hamming Distance
iji-j / $1/4$
iji-j $i\rightarrow j$ $1/4$
$i\rightarrow j$ iji-j $1/5$
Table 5: SHD penalties for various discrepancies between ground truth and predicted edges. $i\rightarrow j$’: oriented edge, ’iji-j’: unoriented edge, ’//’: no edge
@@ G.1 Label accuracy
Table 6: Label accuracy (%). Task concepts are as follows: dysp (Asia), Akt (Sachs), BP (Alarm), PropCost (Insurance), R5Fcst (Hailfinder), parity (cMNIST), mouth slightly open (CelebA), Survival (CUBC), pneumothorax (Pneumoth.). Uncertainties represent 22 sample mean $\sigma$ across 5 runs.
@@ G.2 Task accuracy using true graph
Table 7: Task accuracy (%) using the true and predicted graph. Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder). Uncertainties represent 22 sample mean $\sigma$ across 5 runs.
@@
C2BM True graph $71.0_{\pm 2.3}$ $5.0_{\pm 1.3} 66$$6.4_{\pm 1.8}$ $62.1_{\pm 1.5}$ $73.0_{\pm 1.6}$
Predicted graph $71.4_{\pm 1.7}$ $5.3_{\pm 1.1} 66$$6.4_{\pm 1.5}$ $62.5_{\pm 1.4}$ $74.1_{\pm 1.8}$
@@
G.3 Ablation study on Causal Discovery
In this section, we evaluate the sensitivity of the causal discovery component in our pipeline to the choice of method for constructing a causal graph. To do so, we compare our method of choice for causal discovery, i.e., Greedy Equivalence Search (GES) (chickering2002), with different widely-used causal discovery algorithms that recover a CPDAG. Each of these methods is evaluated with and without refinement via the retrieval-augmented generation (RAG)-enhanced language model (LLM) employed in our study. Furthermore, we compare all these methods against the use of the LLM alone, as well as our retrieval-augmented LLM (LLM + RAG) applied directly to discover the whole graph. Specifically, the LLM is prompted with the causal prompt shown in App. C.3, with additional retrieved context appended to the query in the LLM+RAG setting.
@@ G.3 Ablation study on Causal Discovery
  • GES (score-based): The algorithm selected in our study. A detailed description is provided in App. C;
  • PC Algorithm (spirtes2000causation) (independence-based): A classical independence-based method that first estimates the undirected structure of the causal relations through a sequence of conditional independence tests, then orients edges using a set of predefined orientation rules (colombo2014). In our implementation, we use the PC algorithm provided by the causal-learn library (zheng2024), with a chi-squared independence test and a significance level of 0.05;
  • Fast GES (ramsey2017) (FGES) (score-based): A computationally efficient variant of GES that improves performance by storing intermediate evaluations and parallelizing expensive operations, enabling its application to large datasets (andrews2019; ramsey2017). In our implementation, we use the FGES algorithm provided by the py-tetrad library (ramsey2023py), employing the Bayesian Dirichlet Equivalent Uniform (BDeu) score for structure evaluation (heckerman1995learning).
The results are evaluated on all the bnlearn datasets, for which the true causal graphs are known. Results are presented in Tab. 8.
Table 8: Structural Hamming distance and number of mistaken edges between the true and learned causal graphs for each tested method. The maximum number of errors (Max errors) and the average number of errors make by a random classifier (Random Classifier) are also provided for reference.
@@
Number of mistaken Max errors 28 55 351 666 1540
edges Random Classifier 9.33 18.33 117 222 513.33
LLM 13 20 351 543 107
LLM + RAG 12 12 226 Time limit Time limit
PC 6 10 26 13 45
FGES 3 17 23 14 39
GES 3 17 19 13 22
PC + LLM (+ RAG) 6 9 22 9 43
FGES + LLM (+ RAG) 1 8 15 11 36
GES + LLM (+ RAG) 1 7 18 10 22
As shown in the results, GES refined with the RAG-augmented LLM is the overall best-performing method in the majority of cases, both in terms of structural Hamming distance and number of mistaken edges. In instances where it is not the best, it still consistently ranks among the top methods. Notably, both the standard causal discovery method (GES) and the use of the RAG-augmented LLM contribute positively to performance. Due to the substantial computational time required by the LLM+RAG-based causal discovery approach, we excluded experiments on datasets for which execution exceeded 24 hours.
G.4 Ablation study on LLM type
The LLM or the design of the LLM prompt could condition the causal graph refinement. To assess this, we present an ablation study in which we fix the incomplete causal graph generated by the causal discovery algorithm (GES), and later evaluate different LLMs, with different prompts, for the graph refinement step. Specifically, we evaluate the impact of four different prompting strategies with three LLMs (GPT-4o, 200M parameters; and GPT-4o-mini, 8M parameters). The considered strategies are as follows:
  • Minimal prompting: simply asks the LLM to identify causal relations without additional guidance;
  • Instruction prompting: provides a more detailed explanation of what constitutes a causal relation;
  • Few-shot prompting: proposed in ye2024investigating, combines instruction prompting with a few examples;
  • Chain-of-Thought (CoT): proposed in wei2022chain, encourages the model to generate intermediate logical steps before generating the answer. This is the approach we used in the original manuscript.
Tab. 9 compares the number of mistaken edges and the custom Hamming distance between the true and the learned DAG. Causal reliability of standard, flat, CBMs is also reported for reference. The results show that, using GPT-5, causal reliability improves on average w.r.t. GPT-4o, and GPT-4o-mini, especially on the Sachs dataset. Despite moderate variation in the Sachs dataset, the performance is generally very robust to the prompt strategy (made an exception for naive minimal strategy), and causal reliability is largely superior to standard CBMs in all cases.
Table 9: Structural Hamming distance and number of Mistaken Edges between true and learned DAG across datasets, using different LLMs and prompting strategies. Uncertainty represents 22 sample mean $\sigma$ across 3 runs.
Metric LLM Prompting Strategy Asia Sachs Insurance Alarm Hailfinder
[…]
G.5 Ablation study on RAG
@@
Mistaken No context 12 10
edges ratio RAG context 7 10
@@
G.6 Sample complexity of graph discovery
We study the effect of dataset size and graph size on the quality of C2BM’s graph construction pipeline. We explored this with a sensitivity study, running the causal graph pipeline (causal discovery with GES + LLM refinement with CoT prompt) across all datasets with an available ground-truth graph and comparing the number of mistaken edges between the true and the learned DAG. For each dataset, we varied the number of data points $N$ from $100$ to $10000$.
Table 11: Number of mistaken edges between the true DAG and the learned graph using the C2BM’s graph construction pipeline, evaluated across datasets and data sizes. For reference, the reliability of standard flat CBMs is also reported.
Dataset Flat CBMs C2BM’s causal graph
Data size ($N$) $\rightarrow$ 100 500 1000 5000 10000 (Paper)
Asia 11 6 3 1 1 1
Sachs 23 15 12 11 7 7
Insurance 74 47 35 24 22 18
Alarm 78 31 16 9 9 9
Hailfinder 117 56 47 44 22 22
Results are presented in Tab. 11.
A few considerations emerge:
  • C2BM’s causal graph is consistently more causally reliable than the flat structure implicitly assumed by standard CBMs, regardless of the dataset size.
  • As expected, increasing the number of data points leads to better alignment between the estimated and true causal graphs. Causal reliability tends to remain stable at larger data sizes.
  • We observed that the data size threshold for reliable performance does not strictly depend on graph size. We speculate this is due to the varying impact of LLM-based refinement across datasets. In some cases, an effective background knowledge can compensate for limited data.
G.7 Sensitivity to graph corruption
In this section, we empirically assess the robustness of C2BM to graph misspecification and corruption. Although C2BM is theoretically a universal approximator for the final prediction task, independent of the specific causal graph (see Appendix D.1), we complement this result with an empirical validation.
  • Adversarial Graph Corruptions. We first evaluate robustness by altering a percentage $p$ of graph edges, chosen randomly, with one of the following operations: edge flipping, addition, or removal. The resulting performance across datasets and corruption levels is reported in Tab. 12.
  • Progressive Flattening into Standard CBMs. As a second evaluation, we progressively transform the graph into the flat structure assumed by standard CBMs, by connecting a percentage $p$ of nodes directly to the prediction task while removing their outgoing edges. Results are shown in Tab. 13.
Table 12: Task accuracy (%) under edge-level adversarial corruptions. A percentage $p$ of edges is altered through flipping, addition, or removal. Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder). Uncertainties represent 22 sample mean $\sigma$ across 3 runs.
Dataset / pp 0.05 0.1 0.2 0.4 0.6 0.8 1.0
Asia 71.8±0.6 71.6±0.5 71.4±0.6 71.7±1.3 70.2±1.9 71.4±1.2 70.2±0.9
Sachs 65.1±1.9 65.6±2.0 64.9±2.0 64.6±1.2 65.5±1.5 65.2±1.5 65.0±2.0
Insurance 67.2±1.6 67.0±3.1 67.5±2.3 66.1±1.1 67.0±2.6 66.8±2.6 67.2±3.1
Alarm 61.9±2.6 61.7±2.1 61.5±2.5 61.9±2.7 60.6±1.9 60.9±1.6 61.4±1.7
Hailfinder 74.0±1.5 73.1±3.0 72.3±2.2 72.3±3.2 73.0±2.1 71.8±1.6 72.4±1.4
Table 13: Task accuracy (%) under progressive graph flattening into standard CBMs. A percentage $p$ of nodes is directly connected to the task output. Task concepts are as follows: dysp (Asia), Akt (Sachs), PropCost (Insurance), BP (Alarm), R5Fcst (Hailfinder). Uncertainties represent 22 sample mean $\sigma$ across 3 runs.
Dataset / pp 0.05 0.1 0.2 0.4 0.6 0.8 1.0
Asia 72.0±1.0 72.0±1.0 71.5±0.3 71.7±0.8 71.7±0.5 71.2±1.6 71.1±1.1
Sachs 65.5±2.2 64.8±3.0 65.8±1.6 64.7±1.5 65.2±1.2 65.1±2.7 64.9±2.2
Insurance 67.4±2.6 66.2±1.7 66.3±2.5 65.6±3.3 65.3±2.4 66.9±1.8 64.7±3.1
Alarm 62.6±2.9 61.6±2.6 61.6±2.0 60.5±2.6 61.2±1.8 60.6±3.2 61.3±1.9
Hailfinder 73.1±1.1 73.2±1.9 72.5±0.9 72.7±2.9 73.0±2.6 73.0±1.7 72.9±1.9
Across both corruption strategies, we find that task accuracy remains stable, even when the causal graph is heavily perturbed. These empirical results support the theoretical claim that C2BM is robust to graph misspecification.
G.8 Effect of single-concept interventions
G.9 Decomposing interventional accuracy
@@
Figure 12: Task accuracy improvement (%) in predicting downstream variables after intervening on groups of concepts up to progressively deeper levels in the graph hierarchy. The metric is averaged across all downstream variables (Left), the task only (Middle), and all concepts (Right) Total label accuracy. Uncertainties represent 22 sample mean $\sigma$ across 5 runs.